OpenStudy (anonymous):

so if i have points (1,4) and (-1,1) how do i get it in standard form

OpenStudy (anonymous):

um i know how i'm just too tired to tell you i'm sorry

OpenStudy (ranga):

Find the slope of the straight line from the given two points: m = (y2-y1) / (x2 - x1) The equation of a line knowing the slope and any point is: y - y1 = m(x - x1) Then rearrange it to standard form to look lie: Ax + By = C

OpenStudy (anonymous):

so how would i arrange it if i got y=4 and 3/4 x -3/4 ?

OpenStudy (ranga):

Slope m = (4 - 1) / (1 - (-1)) = 3/2 Equation of the line with slope 3/2 passing through (1,4) is: y - 4 = 3/2(x - 1) multiply both sides by 2 to get rid of the fraction 2y - 8 = 3x - 3 add -2y to both sides -8 = 3x - 2y - 3 add 3 to both sides -5 = 3x - 2y or 3x - 2y = -5 is the standard form of the line. Just to confirm put the two given points into this equation and make sure it satisfies the equation.

OpenStudy (ranga):

Also, when coming up with standard form Ax + By = C make sure that A > 0 If not multiply throughout by -1 to make A positive.

OpenStudy (anonymous):

thanks